Fermi National Laboratory
 

Accelerator Update for November 17, 2006 - November 20, 2006

 

Friday November 17

The day shift began with the Tevatron (TeV) in standby while other machines recovered from the power glitch, with Operations monitoring a Recycler stash (68.9E10), with the Antiproton Source (Pbar) stacking (50.3E10), with Linac, Booster, the Main Injector (MI), NuMI, and MiniBooNE recovering, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access.

The Recycler completed a stash of antiprotons from Pbar at 9:47 AM.

Operations prepared the TeV transfer hall and sectors A-E for access at 10:08 AM.

At 10:48 AM, mechanical technicians reported finding a large ball of ice and a frozen water line at TeV sector C4. They began melting the ice.

Operations established beam to MiniBooNE at 10:49 AM.

The Recycler completed a stash at 1:16 PM.

Operations established beam to NuMI at 1:23 PM.

D0 came out of access at 1:34 PM.

Operations prepared the TeV F sector for an access at 1:37 PM to investigate a sump problem.

Operations put the Booster into dump mode at 1:49 PM. Booster experts began a study period.

Linac experts switched the Preaccelerator from the H- Source to the I- Source at 1:50 PM.

At 2:56 PM, FESS techs reported that the E35 sump pump had failed. Replacing the pump would take about nine hours, so they will install a portable pump. This work took about an hour.

Operations established beam to MiniBooNE at 3:44 PM.

Operations put the TeV through a dry squeeze at 9:33 PM.

At 10:08 PM, Operations reported having a problem with stochastic cooling for the Recycler. An expert came in and investigated.

At 11:21 PM, Operations reported a problem with the Pbar target air-pressure. The pump tripped off and wouldn’t restart. Stacking was stopped. Operations contacted an expert.

Operations established a 14x0 study-store at 11:31 PM (14 bunches of protons and zero antiprotons).

Saturday November 18

The midnight shift began with Operations monitoring a stash (156.6E10) and stack (108.7E10), with Pbar not stacking, and with MiniBooNE and NuMI taking for beam.

Operations put the TeV into shot setup at 12:48 AM.

Recycler experts reported at 2:22 AM that they had completed their work on the vertical dampers.

A Pbar expert reported at 3:56 AM that the air blower had a broken shaft and it was replaced.

Operations established store 5080 at 4:02 AM with an initial luminosity of 105.8E30.

The Recycler completed a stash at 4:23 AM.

Pbar resumed stacking at 4:24 AM.

The Recycler completed a stash at 7:19 AM.

At 9:50 AM, Operations reported problems with Recycler cooling. An expert recommended turning off the dampers, which was done.

At 10:04 AM, the I- Source tripped off. It reset okay.

Operations reported at 10:23 AM that turning off the dampers had helped the Recycler emittances.

At 11:19 AM, Operations had put the Recycler into shot setup, but Accumulator problems stopped the transfer. Operations put Pbar back into stacking mode at 11:34 AM.

The Recycler completed a stash at 12:20 PM. Operators had found and restored a bad file.

The Recycler completed a stash at 3:55 PM.

Operations terminated store 5080 at 6:04 PM.

Operations had D0 and CDF prepared for an access by 6:27 PM

At 7:17 PM, Operations had prepared the Recycler to stash. The first transfer didn’t make it in. Operators investigated and found that the kicker timing was off.

The Recycler completed a stash at 7:49 PM.

CDF and D0 were out of access by 8:38 PM.

Operations put the TeV through a dry squeeze at 9:12 PM.

Operations put the TeV into shot setup at 9:43 PM.

At 10:49 PM, a Booster kicker tripped off on a fan fault and would not reset. Operators bypassed the kicker and called in an expert.

The Booster expert had the kicker fan replaced by 11:56 PM.

Sunday November 19

The midnight shift began with the TeV in shot setup, with Operations monitoring a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 5082 at 1:42 AM with an initial luminosity of 181.2E30.

The Recycler completed a stash at 1:50 AM.

The Recycler had completed two more stashes by 7:03 AM.

Recycler experts began a study period at 2:24 PM.

The Recycler had completed three more stashes by 5:17 PM.

At 9:11 PM, Operations reported having problems with the Pelletron. Operations called in an expert.

The Recycler expert had the Pelletron back on by 10:18 PM.

The Recycler completed a stash at 11:27 PM.

Operations terminated store 5082 at 11:34 PM.

Operations prepared CDF for an access at 11:43 PM.

Monday November 20

The midnight shift began with Operations monitoring a stash, with Pbar stacking, with CDF in access, and with NuMI and MiniBooNE taking beam.

CDF came out of access at 1:06 AM.

The Recycler completed a stash at 1:41 AM.

Operations put the TeV through a dry squeeze at 1:42 AM.

Operations put the TeV into shot setup at 2:08 AM.

Operations established store 5084 at 4:18 AM with an initial luminosity of 210.12E30. (This was the third largest luminosity ever established in the TeV.)

The Neutron Therapy Facility reported at 7:26 AM that they would be treating three patients this morning.

The Recycler completed two more stashes by 7:53 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI to conduct some studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/20/2006    email Fermilab